OpenStudy (anonymous):

Show that for t>0 logt is not a polynomial.

OpenStudy (amistre64):

umm.. it doesnt fit the definition of a poly maybe?

OpenStudy (anonymous):

what is the definition of log(t)?

OpenStudy (amistre64):

gotta ask a lumberjack hah!

OpenStudy (anonymous):

has only one zero? hard to know what this question is asking

OpenStudy (anonymous):

amistre i have a question for you. solve for x: 3x+1

OpenStudy (anonymous):

u have o prove that its not a polynomial...thats what given in my book's exercise..

OpenStudy (amistre64):

they always say proving a negative cant be done

OpenStudy (anonymous):

still not sure what it is asking. you need some definition to work with. presumably you are supposed to use some property of the logs that polynomials don't have, but there are so many

OpenStudy (amistre64):

arent logs exponents?

OpenStudy (anonymous):

May be a polynomial is something which has only terms like ax^b

OpenStudy (anonymous):

ln(ab)=ln(a)+ln(b) for example. what polynomial has that property?

OpenStudy (anonymous):

u have to prove it rigourously..

OpenStudy (amistre64):

define rigorous lol

OpenStudy (anonymous):

well you cannot prove it rigorously without appealing to a definition is my guess.

OpenStudy (anonymous):

means the math teacher likes it.

OpenStudy (anonymous):

Yes, you can't say that an apple is red, or is not orange, without defining what red and orange are!

OpenStudy (anonymous):

no polynomial has a vertical asymptote. no polynomial is undefined at 0.

OpenStudy (anonymous):

i said t>0 satellite73

OpenStudy (anonymous):

so? still any asymptote yes? limit as x->0 ln(x) = - infinity

OpenStudy (anonymous):

use this definition:In mathematics, a polynomial is an expression of finite length constructed from variables (also known as indeterminates) and constants, using only the operations of addition, subtraction, multiplication, and non-negative integer exponents.

OpenStudy (anonymous):

no poly has that property.

OpenStudy (anonymous):

prove that no polynomial has that property..

OpenStudy (anonymous):

???

OpenStudy (anonymous):

well its easy

OpenStudy (anonymous):

k. accepted..but would like noncalculus proof..

OpenStudy (anonymous):

the more i think about this the less sense it makes. i am not really sure what it is after. probably something along the lines of "sin(x) is not a polynomial because a polynomial of degree n has at most n real zeros"

OpenStudy (anonymous):

so probably some property of the logs that polynomials do not possess. already suggested ln(ab)=ln(a)+ln(b)

OpenStudy (anonymous):

well deriving from the main definition is the job..

OpenStudy (anonymous):

the book told so..

OpenStudy (anonymous):

i think all the properties will come from the definition.

OpenStudy (anonymous):

but f(xy)=f(x)+f(y) may be true for some polynomials..

OpenStudy (anonymous):

so tat property cant be used!!!

OpenStudy (anonymous):

f(xy)=f(x)+f(y) is not true for any polynomial.

OpenStudy (anonymous):

well f(2x)=constant+f(x) does not hold for any polynomial but with log

OpenStudy (watchmath):

With calculus: Suppose it is equal to some \(p(t)\) then \(\lim_{t\to \infty} \log t/p(t)=1\) On the other hand, by L'hospital rule \(\lim \log t/p(t)=0\)

OpenStudy (anonymous):

@amogh what about the zero polynomials?

OpenStudy (anonymous):

what is with the latex?

OpenStudy (anonymous):

its screwed up...

OpenStudy (anonymous):

I don't think constants come into polynomials, do they?